[36205] in Hotline Meeting
Requested Filter Changes - Bldg. 9 - 12/19 & 20/96
daemon@ATHENA.MIT.EDU (OPCTEMP2%MITPPL.bitnet@mitvma.mit.)
Mon Dec 16 10:27:49 1996
Date: Mon, 16 Dec 1996 10:26 EDT
From: OPCTEMP2%MITPPL.bitnet@mitvma.mit.edu
To: shutdowns@MIT.EDU, jobrien@MIT.EDU, westlunc@MIT.EDU, pcollymo@MIT.EDU,
roneil@MIT.EDU
X-Envelope-To: jobrien@mit.edu, pcollymo@mit.edu, roneil@mit.edu,
shutdowns@mit.edu, westlunc@mit.edu
X-Vms-To:
IN%"shutdowns@mit.edu, jobrien@mit.edu, westlunc@mit.edu, pcollymo@mit.edu, ron
eil@mit.edu"
The following requested filter changes for Bldg. 9 for 12/19 and 20 are
for your information and comment. Would you please let me know if this
has your approval or not.
Thank you.
Andrea Nichols
Operations Center/PM Office
8-7872
Building: 9 - 551 PA, 550 PA, 550A, 451RA, 351AV, 255 CR, 0050 MR
and 045.
System: Fan coil units (will remain on)
Date: 12/19 and 20/96
Day: Thursday and Friday
Time: 7:00 AM - 2:00 PM each day
Duration: 2 days
Reason: 51 filter changes on fan coil units
Effect: None - each fan coil unit will remain operational.
Filter people will come in, remove old filter and replace with
the new one.